OpenStudy (anonymous):

please help one math problem A construction crew is lengthening a road. The road started with a length of 53 miles, and the crew is adding 3 miles to the road each day. Let L represent the total length of the road (in miles), and let D represent the number of days the crew has worked. Write an equation relating L to D . Then use this equation to find the total length of the road after the crew has worked 38 days.

OpenStudy (anonymous):

L=53+3D "L" is the total length "3D" represents 3 miles times the # of days worked

OpenStudy (anonymous):

Questions?

OpenStudy (anonymous):

that is the equation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its saying: the total length equals 53 miles plus 3 miles times however many days worked

OpenStudy (anonymous):

so length is 167

OpenStudy (anonymous):

eq 53+3(D) = L

OpenStudy (anonymous):

yes 53+3(38)= L so L= 167

OpenStudy (anonymous):

which one do i use this one 53+3(D) = L or yours?

OpenStudy (anonymous):

both are the same. you just rearranged the equation

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

ur welcome